# Semantics-Empowered Space-Air-Ground-Sea Integrated Network: New Paradigm, Frameworks, and Challenges

### v1

@@ IV-D Lessons Learned from This Section
This section provides a detailed review of the meaning-oriented semantic communication system. The metrics reflecting the “meaning similarity” between the original and reconstructed data serve as key guidelines for the system design, and it is worth noting that the accurate representation of the “meaning similarity” for the utilized evaluation metrics is the basis to improve the superior system performance. As the fundamental method for system design, DNNs are considered as the absolute core of the meaning-oriented communication system. When designing the system, considering the limited device capabilities in the SAGSIN, incorporating meaning-enhancement techniques into DNNs is a powerful approach to further improve the system performance. As a summary, this section provides an idea for designing the meaning-oriented semantic communication system: starting from maximum meaning similarity metrics, establishing DNN-related models, and then proposing targeted meaning-enhancement techniques.
@@
Before the review, we firstly elaborate on the reason for introduction of task-oriented communications. Traditionally, both technical communications and meaning-oriented semantic communications adopt separate reconstructor and decision-maker design, where the receiver are expected to firstly reconstruct the estimation of source data by inverse process of transmitter, and secondly generate task decision results based on the estimation of data. The reason why reconstruction process is imperative is that the receiver cannot distinguish which part of received data is task-oriented, and thus all the source data must be firstly reconstructed for further decision. In stark comparison, task-oriented communications jointly design the reconstructor and decision-maker, where an intelligent module at the receiver implicitly processes the received data and directly produces task decision results. This joint design has a salient characteristic of non-symmetry, since no explicit reconstruction process corresponding to pre-processing (e.g., encoding or modulating) occurs at the receiver. Joint reconstructor & decision-maker design is feasible for receiver side to generate moderate decision results, because the received data contain all the information that is needed for tasks, and thus reconstruction is unnecessary for decision process except when the task itself is to reconstruct source data. Moreover, such non-symmetric design of joint reconstructor & decision-maker module is proven to be more simpler in computation than separate design by state-of-the-art literature to be elaborated in Section V-B, which perfectly addresses the issue of limited computing resources in the SAGSIN. On the other hand, since the goal of task-oriented communications is just yielding perfect decision at receiver side for task actuation at the terminal, semantic extractor at the transmitter can be further refined to be also task-oriented. Specifically, the semantic extractor introduced in Section IV can extract all of the semantic information contained by source data. However, the extracted semantic information is not all useful for task decision. Therefore, a more intelligent task-oriented semantic extractor can be introduced, which extracts only effectiveness-related information (i.e., task-related information) of source data to generate effectiveness-related semantic representation for further transmission. Since the decision-maker at the receiver needs only task-related information for decision, this effectiveness-related information extractor is feasible for task-oriented communications. Furthermore, the amount of data gets thoroughly declined as compared to extracting all the semantic information, which also tackles the challenge of limited communication capabilities in the SAGSIN.
@@
In task-oriented communications, a radical departure of communication goal from reconstruction to enhancing ultimate effectiveness represents that the actuation process at the terminal along with abovementioned perception, communication, and computing processes is taken into account in network design, by which a new PCCAIP is developed as demonstrated in Fig. 11. Similar to Section IV, we assume that the raw data have been sampled by intelligent sampler. Specifically, in order to reduce the redundancy caused by effectiveness-unrelated information, we utilize an effectiveness-related information extractor as semantic extractor at the transmitter to remove unrelated information and generate effectiveness-related representation, which will be further coded and modulated by JSCC encoder and modulator to generate transmitted message. At the receiver, different from Section III and Section IV where received messages are reconstructed or recovered by demodulator, decoder, and reconstructor, only one intelligent DL-based actuator receives messages as input and outputs directly the task decision results. Similarly, edge servers and cloud data centers provide knowledge bases and support the distributed learning process of the whole network. By intelligent actuator, the actuation of task will affect the physical world at the transmitter via feedback and effectiveness is thus yielded, which is a closed-loop framework. In the following subsections, we firstly introduce metrics to measure the ultimate effectiveness of task-oriented communications; then, we detail the EYTs for the task-oriented semantic communication system according to specific task categories, including image processing services, remote control services, and other computing-intensive services.
@@ V-A1 Accuracy of tasks
The tasks related to classification, such as image/video classification, image retrieval, and audio recognition, necessitates the correctness or accuracy of output results. For these classification tasks, the samples can be coarsely divided into four categories, which are true positive (TP), true negative (TN), false positive (FP), and false negative (FN). Denote the number of classified samples as , and a specific category of samples as , , , or , respectively. Generally, the precision rate is defined as
@@
where is the number of total samples. The classification accuracy defined in equation (31) describes how many samples of one category is correctly classified, which is also called top-1 accuracy. As a general case of top-1 accuracy, top- accuracy is defined as the ratio of samples whose real category (also called ground truth) belongs to the categories with highest degree of confidence.444The degree of confidence is usually the direct output of DL-based classifier, which represents the probabilities of all the categories. In the cases where the number of categories is small, top-1 accuracy is usually adopted. When the number of categories is large, in order to avoid feature overlapping among categories (a sample has similar degree of confidence on several categories), top- accuracy is more common. In specific classification tasks where the negative impact of FPs is comparable to that of TNs, is sufficient for serving as ultimate effectiveness metric. However, in the SAGSIN scenario, the cost yielded by different classification results may be remarkably distinct, implying that the impact of FPs and TNs should be considered as different in more general situations. By adopting this idea, the F-measure metric is proposed[178] to endow different weight for FPs and TNs by a hyper parameter , defined as
@@
where is the number of categories, is the ratio of category among all samples, and is the F1-score of category calculated by (33). can serve as an ultimate effectiveness metric for multi-classification tasks. For object detection and image segmentation tasks, we are interested in specific classified results in the pixel level. That is, the accuracy of these tasks are determined by the ratio of correctly classified pixels in each image, i.e., whether the detected area or segmented area overlaps the ground truth area. The intersection over union (IoU) metric is defined as the ratio of the intersected area by the union area between the task output results and the ground truth , that is
where represents the area scaled by the number of pixels. It is worth noting that the ground truth is usually labeled manually, which represents the perception of human on the task results. Because these detection and segmentation tasks necessitate evaluation from human during task decision, the IoU between classification and real results (i.e., approximate perception of human) is competent for these tasks.
@@ V-A2 Preciseness of tasks
The tasks related to operation and control, such as automatic driving and UAV route tracking, are subject to the preciseness of the commands generated by the decision-maker, for only the precise and exact commands based on the environment and current situation will yield positive effect on the tasks, while inappropriate commands will hinder the effectiveness of tasks. For these operation tasks, the most determinant factor is system stability, which means that the generated commands must render the state of the system convergent. Here we take wireless network control system (WNCS) as an example to introduce the preciseness metrics that measure the ultimate effectiveness. Consider a discrete-time process denoted by which is controlled by command and interrupted by a noise normally distributed with zero mean and covariance matrix . The process is assumed as a linear time-invariant system, that is[179, 180]
where are the state transition matrix and control coefficient matrix, respectively. One of the most common preciseness metric is the MSE of process , with a tacit assumption that the goal of control task is maintaining as close as zero. Usually we utilize the long-term average MSE as preciseness metric of the WNCS, denoted by
@@ V-B Techniques: Towards Effectiveness Yielding
In the task-oriented semantic communication system, the extraction, dissemination, and ultimate actuation is remarkably distinct from the meaning-oriented semantic communication system discussed in Section IV. Task-oriented semantic communication systems are designed towards task actuation at the terminal, and thus the resource consumption of perception, communication, and computing get reduced, which satisfy the resource-efficient demand in the SAGSIN. In this subsection, we review the EYTs applied in the task-oriented communications which facilitate typical services in the SAGSIN, such as image processing, remote control, and other computing-intensive services. A summary of these EYTs can be seen in Table V.
@@
Accuracy, MSE | Multi-device scheduling |
@@ V-B1 EYTs for image processing services
Image processing tasks include image classification, objective detection, image segmentation (semantic segmentation included), and image retrieval, etc, and the main differences among these tasks are illustrated in Fig. 12. These services require only for accurate results at the terminal instead of exact restoration of image itself. Therefore, a joint transmitter-actuator training method can perfectly and efficiently conduct these image processing tasks in the SAGSIN with limited computing resources, benefiting from lite model size and high computing speed. Image classification services. In [181], the authors study a task-oriented image classification system from a freshness perspective. In this system, the transmitter encodes the images as analog code symbols (i.e., code symbols in a float type) for transmission, and the intelligent classifier deployed at the receiver directly yields the category of a specific image based on received symbols without reconstruction of image. Taking the age of task information (AoTI) as objective function, the optimal arrival rate and analog code length are respectively solved to achieve highest classification accuracy. The authors in [182] theoretically analyze a task-oriented communication system for classification based on robust information bottleneck theory, by which both high relevance of information on tasks and low data distortion are guaranteed, achieving high inference accuracy on images. An adaptive compression method of the task-oriented communication system for classification from a resource allocation perspective is proposed in [183]. According to the relevance on tasks, the compression ratio and resource consumption are jointly optimized by a two-stage iteration algorithm. Simulation results demonstrate that the proposed adaptive compression scheme saves 80% transmission data while maintaining the classification accuracy, implying that both high ultimate effectiveness and high efficiency is achieved. For the sake of realization of task-oriented communications in the SAGSIN scenarios based on digital modulation, [184] proposes a possible solution which utilizes a spiking-neural-network-based (SNN-based) semantic communication system for image classification task. Compared with other NNs such as CNN, RNN which have been introduced in Section IV-B, a remarkable feature of SNN [201] is that the output of last layer is only a vector consisting of binary elements, i.e., , where is the number of neurons at the last layer of SNN. This remarkable feature implies that SNN directly outputs digital symbols that can be modulated by constellations such as BPSK without quantization procedure, reducing distortion of effectiveness-related information caused by extra quantization. [184] considers a task-oriented semantic codec equipped with SNN over the BSC or BEC, and provides simulation results on the classification accuracy with regards to transition/erasure probability. To enhance the robustness of task-oriented communication system against adversarial attacks, the authors in [185, 186] focus on anti-semantic-noise semantic communications for robust image classification. The GAN is utilized for generating image data set polluted by semantic noise which cannot be distinguished by human but may mislead the intelligent classifier to yield wrong category results. The authors introduce mask auto-encoder to encounter the negative effect of semantic noise, and propose a semantic codebook method similar to [202, 203] in order to extract only the most significant features (i.e., the most task-relevant semantics) for transmission. By these means the classification accuracy does not decline because of semantic noise compared with noise-free scenarios.
@@
Objective detection services. [187] considers task-oriented edge video transmission and detection. Specifically, a temporal entropy model and a temporal-spatial integrated model are respectively proposed to extract effectiveness-related information from both current frame and previous frame, according that video sources are temporally continuous and thus the information involved by previous frame can be neglected for feature extraction in the current frame. The ultimate effectiveness in specific tasks, e.g., precision (based on IoU thresholds) of pedestrian distribution detection and multi-device joint objective detection, gets enhanced by task-oriented model design while achieving better rate-effectiveness trade off. Image segmentation services. [188] studies Internet-of-vehicles-based (IoV-based) task-oriented communication networks for real-time image segmentation. The cars in the front row capture, process, and transmit environment images in real time by swin Transformer encoder, and the cars in the back row conduct image segmentation tasks for received images. The intelligent task-oriented transmitter-actuator pair achieves higher mIoU performance than traditional data-oriented codec pair, showing the superiority of ultimate effectiveness of task-oriented design. Image retrieval services. The authors in [189] consider image retrieval issue at the edge, where one edge device captures and transmits images to the edge server, and the server conduct retrieval task to predict the identification of the images. Specifically, two task-oriented schemes namely SSCC and JSCC schemes are utilized for effectiveness-oriented feature extraction, coding, and modulation. Moreover, an fully-connected-network-based classifier at receiver directly conduct retrieval tasks without image reconstruction. Numerical results show that JSCC scheme achieves better task accuracy in a higher actuation speed. Further considering the multi-device scenario at the edge, [190] studies collaborative image retrieval problem based on edge computing. Taking two-device case as an example, edge devices capture images and conduct JSCC-based effectiveness-oriented feature extraction independently, and the JSCC outputs (i.e.,effectiveness-related information representation) are transmitted over a shared multiple access channel using NOMA or OMA modulation. The task actuator at the edge server conducts joint image retrieval task by simultaneously processing information from two edge devices. Such collaborative task actuation achieves performance gain in total retrieval accuracy as compared to separate task actuation design.

### v3

@@ IV-D Lessons Learned from This Section
This section provides a detailed review of the meaning-oriented semantic communication system. The metrics reflecting the “meaning similarity” between the original and reconstructed data serve as key guidelines for the system design, and it is worth noting that the accurate representation of the “meaning similarity” for the utilized evaluation metrics is the basis to improve the system performance. As the fundamental method for system design, DNNs are considered as the absolute core of the meaning-oriented communication system. When designing the system, considering the characteristics of SAGSIN such as highly dynamic channels and limited device capabilities, incorporating meaning-enhancement techniques into DNNs is a powerful approach to further improve the system performance. As a summary, this section provides an idea for designing the meaning-oriented semantic communication system: starting from maximum meaning similarity metrics, establishing DNN-related models, and then proposing targeted meaning-enhancement techniques.
It is worth noting that, the NNs introduced in the meaning-oriented semantic communications will not deteriorate the processing delay performance despite a high computing complexity. Since most of the meaning-oriented semantic communication systems are trained offline, i.e., pre-trained at edge/cloud servers before implementation, the high computation costs in the training phases can be neglected in the communication phases. Moreover, some existing works have quantitatively demonstrated that the processing delay of DL-based semantic communications is shorter than traditional communications. For instance, in [178], the authors do extensive experiments and simulation results show that the image processing delay under DeepJSCC method is only 18 ms in average, while this value under traditional joint photographic experts group (JPEG) source coding and LDPC channel coding method is larger than 30 ms. Therefore, DL-based meaning-oriented semantic communications will be a promising option in implementing task-critical applications requiring for lower delay.
@@
Before the review, we firstly elaborate on the reason for introduction of task-oriented communications. Traditionally, both technical communications and meaning-oriented semantic communications adopt separate reconstructor and decision-maker design, where the receiver are expected to firstly reconstruct the estimation of source data by inverse process of transmitter, and secondly generate task decision results based on the estimation of data. The reason why reconstruction process is imperative is that the receiver cannot distinguish which part of received data is task-oriented, and thus all the source data must be firstly reconstructed for further decision. In stark comparison, task-oriented communications jointly design the reconstructor and decision-maker, where an intelligent module at the receiver implicitly processes the received data and directly produces task decision results.
This joint design has a salient characteristic of non-symmetry, since no explicit reconstruction process corresponding to pre-processing (e.g., encoding or modulating) occurs at the receiver. Joint reconstructor & decision-maker design is feasible for receiver side to generate moderate decision results, because the received data contain all the information that is needed for tasks, and thus reconstruction is unnecessary for decision process except when the task itself is to reconstruct source data. Moreover, such non-symmetric design of joint reconstructor & decision-maker module is proven to be simpler in computation than separate design by state-of-the-art literature to be elaborated in Section V-B, which perfectly addresses the issue of limited computing resources in the SAGSIN.
Moreover, since the goal of task-oriented communications is just yielding perfect decision at receiver side for task actuation at the terminal, semantic extractor at the transmitter can be further refined to be also task-oriented. Specifically, the semantic extractor introduced in Section IV can extract all of the semantic information contained by source data. However, the extracted semantic information is not all useful for task decision. Therefore, a more intelligent task-oriented semantic extractor can be introduced, which extracts only effectiveness-related information (i.e., task-related information) of source data to generate effectiveness-related semantic representation for further transmission. Since the decision-maker at the receiver needs only task-related information for decision, this effectiveness-related information extractor is feasible for task-oriented communications. Furthermore, the amount of data gets thoroughly declined as compared to extracting all the semantic information, which also tackles the challenge of limited communication capabilities in the SAGSIN. Similar to meaning-oriented semantic communication system introduced in Section IV, distributed learning methods are utilized for joint training and knowledge sharing. Through the design of the effectiveness-related information extractor and joint reconstructor & decision-maker, the computing complexity gets significantly decreased compared to the meaning-oriented semantic communication system, resulting in a lightweight distributed learning-based effectiveness-oriented communication system.
@@
In task-oriented communications, a radical departure of communication goal from reconstruction to enhancing ultimate effectiveness represents that the actuation process at the terminal along with abovementioned perception, communication, and computing processes is taken into account in network design, by which a new PCCAIP is developed as demonstrated in Fig. 11. Specifically, in order to reduce the redundancy caused by effectiveness-unrelated information, we utilize an effectiveness-related information extractor as semantic extractor at the transmitter to remove unrelated information and generate effectiveness-related representation. The representation of effectiveness-related information will be further coded and modulated by JSCC encoder and modulator to generate transmitted messages. At the receiver, different from Section III and Section IV where received messages are reconstructed or recovered by demodulator, decoder, and reconstructor, only one intelligent DL-based actuator receives messages as input and outputs directly the task decision results. Similarly, edge servers and cloud data centers provide knowledge bases and support the distributed learning process of the whole network. By intelligent actuator, the actuation of tasks will affect the physical world (e.g., the variation of source data) at the transmitter via terminal execution and effectiveness is thus yielded. The task decision results will be also transmitted back through feedback channel, which will further affect the behavior of transmitter side. Therefore, the effectiveness-oriented semantic communication system is a closed-loop framework.
In the following subsections, we firstly introduce metrics to measure the ultimate effectiveness of task-oriented communications; then, we detail the EYTs for the task-oriented semantic communication system according to specific task categories, including image processing services, remote control services, and other computing-intensive services.
@@ V-A1 Accuracy of tasks
The tasks related to classification, such as image/video classification, image retrieval, and audio recognition, necessitate the correctness or accuracy of output results. For these classification tasks, the samples can be coarsely divided into four categories, which are true positive (TP), true negative (TN), false positive (FP), and false negative (FN). Denote the number of classified samples as , and a specific category of samples as , , , or , respectively. Generally, the precision rate is defined as
@@
where is the number of total samples.
The classification accuracy defined in equation (31) describes how many samples of one category is correctly classified, which is also called top-1 accuracy. As a general case of top-1 accuracy, top- accuracy is defined as the ratio of samples whose real category (also called ground truth) belong to the categories with highest degree of confidence.444The degree of confidence is usually the direct output of DL-based classifier, which represents the probabilities of all the categories. In the cases where the number of categories is small, top-1 accuracy is usually adopted. When the number of categories is large, in order to avoid feature overlapping among categories (a sample has similar degree of confidence on several categories), top- accuracy is more common.
In specific classification tasks where the negative impact of FPs is comparable to that of TNs, is sufficient for serving as ultimate effectiveness metric. However, in the SAGSIN scenario, the costs yielded by different classification results may be remarkably distinct, implying that the impact of FPs and TNs should be considered as different in more general situations. By adopting this idea, the F-measure metric is proposed[191] to endow different weights for FPs and TNs by a hyper parameter , defined as
@@
where is the number of categories, is the ratio of category among all samples, and is the F1-score of category calculated by (33). can serve as an ultimate effectiveness metric for multi-classification tasks.
For object detection and image segmentation tasks, we are interested in specific classified results in the pixel level. That is, the accuracy of these tasks is determined by the ratio of correctly classified pixels in each image, i.e., whether the detected area or segmented area overlaps the ground truth area. The intersection over union (IoU) metric is defined as the ratio of the intersected area by the union area between the task output results and the ground truth , that is
where represents the area (i.e., the number of pixels).
It is worth noting that the ground truth is usually labeled manually, which represents the perception of human on the task results. Because these detection and segmentation tasks necessitate evaluation from human during task decision, the IoU between classification and real results (i.e., approximate perception of human) is competent for these tasks.
@@ V-A2 Preciseness of tasks
The tasks related to operation and control, such as automatic driving and UAV route tracking, are subject to the preciseness of the commands generated by the decision-maker, for only the precise and exact commands based on the environment and current situation will yield positive effect on the tasks, while inappropriate commands will hinder the effectiveness of tasks. For these operation tasks, the most determinant factor is system stability, which means that the generated commands must render the state of the system convergent.
Here we take wireless network control system (WNCS), as illustrated in Fig. 12, as an example to introduce the preciseness metrics that measure the ultimate effectiveness. Consider a discrete-time process denoted by which is controlled by command and interrupted by a noise normally distributed with zero mean and covariance matrix . The process is assumed as a linear time-invariant system, that is[192, 193]
where are the state transition matrix and control coefficient matrix, respectively.
One of the most common preciseness metric is the MSE of process , with a tacit assumption that the goal of the control task is maintaining as close as zero. Usually we utilize the long-term average MSE as preciseness metric of the WNCS, denoted by
@@ V-B Techniques: Towards Effectiveness Yielding
In the task-oriented semantic communication system, the extraction, dissemination, and ultimate actuation processes are remarkably distinct from the meaning-oriented semantic communication system discussed in Section IV. Task-oriented semantic communication systems are designed towards task actuation at the terminal, and thus the resource consumption of perception, communication, and computing gets reduced, which satisfies the resource-efficient demand in the SAGSIN. In this subsection, we review the EYTs applied in the task-oriented communications which facilitate typical services in the SAGSIN, such as image processing, remote control, and other computing-intensive services. A summary of these EYTs can be seen in Table VIII.
@@
Accuracy, throughput, or MSE | Multi-device scheduling |
@@ V-B1 EYTs for image processing services
Image processing tasks include image classification, objective detection, image segmentation (semantic segmentation included), and image retrieval, etc, and the main differences among these tasks are illustrated in Fig. 13. These services require only for accurate results at the terminal instead of exact restoration of image itself. Therefore, a joint transmitter-actuator training method can perfectly and efficiently conduct these image processing tasks in the SAGSIN with limited computing resources, benefiting from lite model size and high computing speed.
@@
Image classification services. In [194], the authors study a task-oriented image classification system from a freshness perspective. In this system, the transmitter encodes the images as analog code symbols (i.e., code symbols in a float type) for transmission, and the intelligent classifier deployed at the receiver directly outputs the category of a specific image based on received symbols without reconstruction of image. Taking the age of task information (AoTI) as objective function, the optimal arrival rate and analog code length are respectively solved to achieve highest classification accuracy. The authors in [195] theoretically analyze a task-oriented communication system for classification based on robust information bottleneck theory, by which both high relevance of information on tasks and low data distortion are guaranteed, achieving high inference accuracy on images. From the resource allocation perspective, an adaptive compression method of the task-oriented communication system for classification is proposed in [196]. According to the relevance on tasks, the compression ratio and resource consumption are jointly optimized by a two-stage iteration algorithm. Simulation results demonstrate that compared to baseline schemes, the proposed adaptive compression scheme saves 80% transmission data while maintaining the classification accuracy, implying that both high ultimate effectiveness and high efficiency is achieved.
For the sake of realization of task-oriented communications in the SAGSIN scenarios based on digital modulation, [197] proposes a possible solution which utilizes a spiking-neural-network-based (SNN-based) semantic communication system for image classification tasks. Compared with other NNs such as CNN, RNN which have been introduced in Section IV-B, a remarkable feature of SNN [214] is that the output of the last layer is only a vector consisting of binary elements, i.e., , where is the number of neurons at the last layer of SNN. This remarkable feature implies that SNN directly outputs digital symbols that can be modulated by constellations such as BPSK without quantization procedure, reducing distortion of effectiveness-related information caused by extra quantization. [197] considers a task-oriented semantic codec equipped with SNN over the BSC or BEC, and provides simulation results on the classification accuracy with regards to transition/erasure probability.
To enhance the robustness of task-oriented communication system against adversarial attacks, the authors in [198, 199] focus on anti-semantic-noise semantic communications for robust image classification. The GAN is utilized for generating image data set polluted by semantic noise which cannot be distinguished by human but may mislead the intelligent classifier to output wrong category results. The authors introduce a mask auto-encoder to encounter the negative effect of semantic noise, and propose a semantic codebook method similar to [215, 216] in order to extract only the most significant features (i.e., the most task-relevant semantics) for transmission. By these means the classification accuracy does not decline because of semantic noise compared with noise-free scenarios.
Objective detection services. [200] considers task-oriented edge video transmission and detection. Specifically, a temporal entropy model and a temporal-spatial integrated model are respectively proposed to extract effectiveness-related information from both current frame and previous frame. The reason why the proposed method is feasible is that video sources are temporally continuous and thus the information involved by the previous frame can be neglected for feature extraction in the current frame. The ultimate effectiveness in specific tasks, e.g., precision (based on IoU thresholds) of pedestrian distribution detection and multi-device joint objective detection, gets enhanced by task-oriented model design while achieving better rate-effectiveness trade-off.
Image segmentation services. [201] studies Internet-of-vehicles-based (IoV-based) task-oriented communication networks for real-time image segmentation. The cars in the front row capture, process, and transmit environment images in real time by swin Transformer encoder, and the cars in the back row conduct image segmentation tasks for received images. The intelligent task-oriented transmitter-actuator pair achieves higher mean IoU performance than traditional data-oriented codec pair, showing the superiority of ultimate effectiveness of task-oriented design.
@@
Image retrieval services. The authors in [202] consider image retrieval issue at the edge, where one edge device captures and transmits images to the edge server, and the server conducts retrieval task to predict the identification of the images. Specifically, two task-oriented schemes namely SSCC and JSCC schemes are utilized for effectiveness-oriented feature extraction, coding, and modulation. Moreover, an fully-connected-network-based classifier at the receiver directly conducts retrieval tasks without image reconstruction. Numerical results show that JSCC scheme achieves better task accuracy in a higher actuation speed. Further considering the multi-device scenario at the edge, [203] studies collaborative image retrieval problem based on edge computing. Taking two-device case as an example, edge devices capture images and conduct JSCC-based effectiveness-oriented feature extraction independently, and the JSCC outputs (i.e., effectiveness-related information representation) are transmitted over a shared multiple access channel using NOMA or OMA modulation. The task actuator at the edge server conducts joint image retrieval task by simultaneously processing information from two edge devices. Such collaborative task actuation achieves performance gain in total retrieval accuracy as compared to separate task actuation design.